OpenStudy (anonymous):

Differentiate the function. y = ln(x^9 sin^2 x)

zepdrix (zepdrix):

Recall:\[\Large \left[\ln(stuff)\right]' \quad=\quad \frac{1}{stuff}(stuff)'\]

zepdrix (zepdrix):

\[\Large y'\quad=\quad \frac{1}{x^9 \sin^2 x}\left(x^9 \sin^2 x\right)'\]

zepdrix (zepdrix):

So we have to take the derivative of that second part there. Product rule will help us out nicely.\[\Large y'\quad=\quad \frac{1}{x^9 \sin^2 x}\left[\color{royalblue}{(x^9)'}\sin^2 x+x^9\color{royalblue}{(\sin^2x)'}\right]\]

zepdrix (zepdrix):

Then just take the derivative of the blue parts, and boom yer done!

OpenStudy (anonymous):

thank you so much!!

OpenStudy (luigi0210):

Wasn't this already answered?